please help me solve this i have no idea of what to do :
0.9(2x+8)=20-(x+5)
thanks!

Found 3 solutions by stanbon, Little girl, MathLover1:
Answer by stanbon(75887)

0.9(2x+8)=20-(x+5)
---
Multiply both sides by 10 to get:
9(2x+8) = 200 - 10(x+5)
18x + 72 = 200 -10x - 50
18x + 72 = 150-10x
28x = 78
x = 39/14
x = 2.7857...
===========
Cheers,
Stan H.
===========

Answer by Little girl(7)

9/10= 0.9
9/10 * (2x+8)= 20-(x+5)
-> 9(2x+8)= 200- 10(x+5)
-> 18x+72= 200-10x-50
-> 28x=78
-> x= 78/28
Hope this can help u ;)

Answer by MathLover1(20849)